### v1

# Zhicong Lin, Feihu Liu, Jiahang Liu§, Jing Liu, and Guoce Xin

Abstract
Banderier, Marchal and Wallner considered Young tableaux with walls, which are similar to standard Young tableaux, except that local decreases are allowed at some walls. In this work, we prove a conjecture of Fuchs and Yu concerning the enumeration of two classes of three-row Young tableaux with walls. Combining with the work by Chang, Fuchs, Liu, Wallner and Yu leads to the verification of a conjecture on tree-child networks proposed by Pons and Batle. This conjecture was regarded as a specific and challenging problem in the Phylogenetics community until it was finally resolved by the present work.
2020 Mathematics subject classification: Primary 05A15; Secondary 05A10; 05A17; 06A07.
Keywords: Young tableaux; Tree-child networks; Exact enumeration; Recursions; Generating functions; Linear extensions.
Contents
1 Introduction
Young tableaux with walls were introduced by Banderier, Marchal and Wallner [1], and their enumeration is quite challenging. The main objective of this paper is to prove an enumerative conjecture (see Theorem 1.1) of Fuchs and Yu [9] concerning two classes of three-row Young tableaux with walls. Tree-child networks are one of the most prominent network classes for modeling evolutionary processes involving reticulation events. Our result together with the work by Chang, Fuchs, Liu, Wallner and Yu [8] leads to the verification of a conjecture due to Pons and Batle [18, Conjecture 1] on counting tree-child networks. Before proceeding to state our main result, we need to introduce some notations and definitions pertaining to Young tableaux with walls.
Let $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ be a partition of $m$, denoted $\lambda\vdash m$. The Young diagram of $\lambda$ is obtained by drawing a collection of left-justified array of $m$ cells, where row $i$ has $\lambda_{i}$ cells reading form bottom to top. For the sake of convenience, the Young diagram of $\lambda$ is also denoted by $\lambda$. A standard Young tableau of $\lambda$ is an assignment of natural numbers $1,2,\ldots,m$ to the cells of the Young diagram, each number appearing once, such that rows are increasing from left to right, and columns are increasing from bottom to top.
Recently, Banderier, Marchal and Wallner [1, 2] considered a variation of standard Young tableaux, called Young tableaux with walls. For a Young diagram, we put a bold red edge (horizontally or vertically) between the consecutive cells which are allowed to be decreasing. Such places where a decrease is allowed (but not compulsory) are drawn by a bold red edge, which we call a “wall”. We refer to this type of Young diagram as Young diagram with walls; see Fig. 1 (a) for an example. As shown in Fig. 1 (b), a Young tableau with walls of $\lambda$ is a filling of the cells by labels $1,2,\dots,m=|\lambda|$, such that each label appears exactly once and the labels are increasing along each row and column (conditions for standard Young tableaux), but two labels separated by a wall need not be increasing.
10681434513151271291116
Figure 1: A Young diagram with walls and a Young tableau with walls.
The enumeration of Young tableaux with walls is a difficult problem. Via bijections, hook-length-like formulas and the density method, some results for the enumeration of rectangular Young tableaux with walls were proved by Banderier, Marchal and Wallner [1]. Banderier and Wallner [2] further considered Young tableaux with periodic walls and reported more enumerative results and bijections with trees, lattice paths or permutations. By employing the kernel method and the lattice path enumeration theory, Liu and Xin [14] obtained the generating function for a class of two-row Young tableaux with periodic walls.
Here we are concerned with two specified classes of three-row Young tableaux with walls. Let $(n,n,k)\vdash 2n+k$ with $n\geq k\geq 0$. We place a vertical wall (bold red edge) between each pair of adjacent cells at the bottom of the Young diagram of $(n,n,k)$. An example is shown in Fig. 2 for $n=7$ and $k=4$. Let $a_{n,k}$ denote the number of all such Young tableaux with walls.
Figure 2: A Young diagram of $(7,7,4)$ with walls at the bottom.
In particular, a closed-form expression for the sequence $(a_{n,n})_{n\geq 0}$ that was registered as [19, A213863] in the OEIS remains unknown. Banderier and Wallner [2, Theorem 4.1] used the density method to derive an asymptotic expression for $a_{n,n}$. Pons and Batle [18, Proposition 7] established the asymptotic behavior of $a_{n,k}$ as $n\rightarrow\infty$ for fixed $k$. On the other hand, it is easy to see that
$a_{n,0}=(2n-1)!!,$ (1)
the odd double factorial, while the sequence $(a_{n,1})_{n\geq 1}$ corresponds to [A122649] in the OEIS [19]. Moreover, by considering the position of the largest label in such a Young tableaux with walls, one can derive a recurrence relation for $a_{n,k}$:
$a_{n,k}=a_{n,k-1}+(2n+k-1)\cdot a_{n-1,k}.$ (2)
The first few values of $a_{n,k}$ are provided in Table 1, for the sake of convenience.
Table 1: Certain values of $a_{n,k}$.
$n,k$ 0 1 2 3 4 5 6
0 1
11 1 1
22 3 7 7
33 15 57 106 106
44 105 561 $15$ 2575 2575
55 945 6555 23220 54120 87595 87595
66 10395 89055 390915 1148595 2462520 3864040 3864040
Next, we introduce the second kind of Young tableaux with walls. Let $(n,n,n)\vdash 3n$ and $n\geq k$. We place a vertical wall (bold red edge) between each pair of adjacent cells at the bottom row of the Young diagram of $(n,n,n)$. By removing any nkn-k bottom cells, we obtain a family of $n\choose k$} deformed Young diagrams with walls. One such diagram is illustrated in Fig. 3 for $n=7$ and $k=4$.
Figure 3: A case in a family of deformed Young diagrams with walls for $n=7$ and $k=4$.
Similarly, a deformed Young tableau with walls of this type is a filling of the cells by labels $1,2,\dots,2n+k$, such that each label appears exactly once and the labels are increasing along each row and column, but two labels separated by a wall need not be increasing. Let $b_{n,k}$ denote the total number of all deformed Young tableaux with walls on this family of deformed Young diagrams with walls. It is well known that
$b_{n,0}=\frac{1}{n+1}\binom{2n}{n}=:\mathrm{Cat}(n),$ (3)
where $\mathrm{Cat}(n)$ is the famous Catalan number appearing as [19, A000108] in the OEIS. The sequence $(b_{n,1})_{n\geq 1}$ was registered as [A000531] in the OEIS [19]. Compared to $a_{n,k}$, there exists no simple recursion for $b_{n,k}$, and computing $b_{n,k}$ appears to be more difficult than computing $a_{n,k}$. A somewhat complicated recurrence relation for a refinement of $b_{n,k}$ was proved in [8, Proposition 3.7]. For the reader’s convenience, we list the first few values of $b_{n,k}$ in Table 2.
Table 2: Certain values of $b_{n,k}$.
$n,k$ 0 1 2 3 4 5 6
0 1
11 1 1
22 2 7 7
33 5 38 106 106
44 14 187 $10$ 2575 2575
55 42 874 7740 36080 87595 87595
66 132 3958 52122 382865 1641680 3864040 3864040
The main result of this paper is the following unexpected relationship between $a_{n,k}$ and $b_{n,k}$, which was originally conjectured by Fuchs and Yu [9].
Theorem 1.1 (Fuchs and Yu’s enumerative conjecture).
For $0\leq k\leq n$, we have
$2^{n-k}\cdot a_{n,k}=(n-k+1)!\cdot b_{n,k}.$ (4)
It is straightforward to verify that
$2^{n}\cdot(2n-1)!!=(n+1)!\cdot\mathrm{Cat}(n),$ (5)
from which we see that Theorem 1.1 holds for $k=0$ in view of (1) and (3). Thus, Theorem 1.1 can be considered as an extension of the relationship (5) between the odd factorial numbers and the Catalan numbers. Theorem 1.1 was inspired by recent work of Chang et al. in [8] and closely connected to a conjecture of Pons and Batle [18, Conjecture 1] on counting tree-child networks that we next elaborate on.
The class of tree-child networks is an increasingly prominent class of phylogenetic networks. Following [18], a phylogenetic network with $n$ leaves is a rooted acyclic digraph without parallel edges such that all nodes belong to one of the following categories:
  • (1)
    A unique root which has indegree 0 and outdegree 1.
  • (2)
    Leaves are nodes that have indegree 1 and outdegree 0. They are bijectively labeled with elements in $\{1,2,\ldots,n\}$.
  • (3)
    Tree nodes which have indegree 1 and outdegree 2.
  • (4)
    Reticulation nodes which have indegree 2 and outdegree 1.
134251342
Figure 4: Two phylogenetic networks: only the right one is tree-child.
A phylogenetic network is tree-child if, for each non-leaf node, at least one of its children is a tree node or leaf. See Fig. 4 for two examples of phylogenetic networks, where only the right one is tree-child. In a phylogenetic network with $n$ leaves, $k$ reticulation nodes and $t$ tree nodes, since the sum of the outdegrees equals that of indegrees, we have
$n+k=t+1.$
Phylogenetic networks and tree-child networks have been extensively investigated from the enumerative aspect [4, 6, 8, 7, 10, 11, 12, 18]. In particular, Pons and Batle [18] proposed the following enumerative conjecture.
Conjecture 1.2 (Pons and Batle [18, Conjecture 1]).
Let $\mathcal{TC}_{n,k}$ be the set of all tree-child networks with $n$ leaves and $k$ reticulation nodes. Then
$|\mathcal{TC}_{n,k}|=\frac{n!}{(n-k)!}\cdot a_{n-1,k}.$ (6)
In fact, $a_{n,k}$ is interpreted in [18] as some Yamanouchi-like words, which are in one-to-one correspondence (see [23, Proposition 7.10.3 (d)]) with the Young tableaux with walls of shape $(n,n,k)$. Conjecture 1.2 stems form the work by Fuchs, Yu and Zhang [11] and has received considerable attention in recent years [8, 10]. In particular, Fuchs, Liu and Yu [10] confirmed Conjecture 1.2 for one-component tree-child networks which are an important building block of (general) tree-child networks (see [7]). Note that several interesting consequences of Conjecture 1.2 have been discussed in [18, 7].
Conjecture 1.2 now follows as a corollary of our main result, Theorem 1.1.
Proof of Conjecture 1.2.
It was proved in [8, Theorem 3.4] that
$|\mathcal{TC}_{n,k}|=\frac{n!}{2^{n-k-1}}\cdot b_{n-1,k},$
which, when combined with Theorem 1.1, implies (6). ∎
The rest of this paper is organized as follows. Sec. 2 is devoted to the proof of Theorem 1.1. Sec. 3 focuses on investigating the generating function for $b_{n,k}$. Via linear extensions of some posets and the principle of inclusion-exclusion, we derive in Sec. 4 a recurrence for $b_{n,k}$. Finally, we end this paper with concluding remarks and related open problems in Sec. 5.
2 Proof of Theorem 1.1
This section is devoted to the proof of Theorem 1.1. Since our proof is somewhat intricate, we first intend to outline the key steps of our proof:
  1. 1.
    Starting from (2), we derive a semi-closed-form expression for $a_{n,k}$ (see Theorem 2.1).
  2. 2.
    We introduce an extension of $b_{n,k}$, denoted $b_{n,m,k}$, by generalizing the deformed Young tableaux with walls. The extension $b_{n,m,k}$ admits a nice recurrence relation (see Proposition 2.5):
    $b_{n,m,k}=(m-k+1)b_{n,m,k-1}+b_{n,m-1,k}+b_{n-1,m,k},$ (7)
    which can be applied to compute $b_{n,k}$ efficiently, as $b_{n,k}=b_{n,n,k}$.
  3. 3.
    Inspired by the favorable recursion for $b_{n,m,k}$ and the conjectured relationship (4) with $a_{n,k}$, we introduce a variant of $b_{n,m,k}$, denoted $\omega_{n,m,k}$, which satisfies a variant of (7) with different initial conditions.
  4. 4.
    To finish the proof, in view of the construction of $\omega_{n,m,k}$ and Theorem 2.1, it remains to show that $\omega_{n,m,k}=b_{n+m,m,k}$. To this end, we establish two key lemmas to verify that $\omega_{n,m,k}$ and $b_{n+m,m,k}$ share the same initial conditions, a part that constitutes the most technical aspect of our proof.
2.1 A semi-closed-form expression for $a_{n,k}$
This subsection aims to prove the following semi-closed-form expression for $a_{n,k}$.
Theorem 2.1.
Let $0\leq k\leq n$. We have
$a_{n,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot(2n+k+i-1)!!,$
where $\gamma_{k}$ are rational numbers satisfying the recurrence relation
$\gamma_{k}=\frac{1}{(3k-3)!!}\left(-\sum_{i=1}^{k}\frac{\gamma_{k-i}}{i!}\cdot(3k+i-3)!!\right)$ (8)
with the initial condition $\gamma_{0}=1$.
Remark 2.2.
We are surprised to learn that Theorem 2.1 was already known in [18, Proposition 6]. However, unaware of their work, we independently discovered this result. Actually, starting from (2) we derive the following recurrence relation for $a_{n,k}$:
$a_{n,k}=a_{n,k-1}+\sum_{i=k}^{n-1}\prod_{j=i}^{n-1}(2(j+1)+k-1)\cdot a_{i,k-1}.$ (9)
Using the above recursion, we produce explicit formulae of $a_{n,k}$ for some fixed small values of $k$ in Proposition A.1, which leads us to find Theorem 2.1.
Banderier and Wallner [2, Theorem 4.1] used the density method to derive an asymptotic expression for $a_{n,n}$. As a special case of Theorem 2.1, we have the following semi-closed exact expression for $a_{n,n}$.
Corollary 2.3.
For $n\geq 0$, we have
$a_{n,n}=\sum_{i=0}^{n}\frac{\gamma_{n-i}}{i!}\cdot(3n+i-1)!!,$
where $\gamma_{j}$ satisfying the recurrence relation (8) and $\gamma_{0}=1$.
Since Theorem 2.1 is crucial in proving Theorem 1.1, we include its proof below, for the sake of completeness.
Proof of Theorem 2.1.
For convenience, we set
$\tilde{a}_{n,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot(2n+k+i-1)!!.$
We aim to show that $\tilde{a}_{n,k}$ and $a_{n,k}$ share the same recurrence relation and initial conditions. Firstly, we have $\tilde{a}_{n,0}=(2n-1)!!=a_{n,0}$. Secondly, the $\tilde{a}_{n,k}$ satisfies the following equation:
$(2n+k-1)\tilde{a}_{n-1,k}=$ $(2n+k-1)\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot(2n+k+i-3)!!$
$=$ $(2n+k+i-1-i)\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot(2n+k+i-3)!!$
$=$ $\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot(2n+k+i-1)!!-\sum_{i=0}^{k}\frac{i\cdot\gamma_{k-i}}{i!}\cdot(2n+k+i-3)!!$
$=$ $\tilde{a}_{n,k}-\sum_{i=1}^{k}\frac{\gamma_{k-i}}{(i-1)!}\cdot(2n+k+i-3)!!$
$=$ $\tilde{a}_{n,k}-\sum_{i=0}^{k-1}\frac{\gamma_{k-i-1}}{i!}\cdot(2n+k+i-2)!!$
$=$ $\tilde{a}_{n,k}-\tilde{a}_{n,k-1}.$
Therefore, we obtain
$\tilde{a}_{n,k}=\tilde{a}_{n,k-1}+(2n+k-1)\tilde{a}_{n-1,k}.$
In view of (2), $\tilde{a}_{n,k}$ and $a_{n,k}$ share the same recurrence relation. Finally, it follows from (8) that
$\tilde{a}_{k-1,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}(3k+i-3)!!=0.$
This is consistent with $a_{k-1,k}=0$, and thus $a_{n,k}=\tilde{a}_{n,k}$. This completes the proof. ∎
2.2 An extension of $b_{n,k}$ and the introduction of its variant $\omega_{n,m,k}$
In this section, we consider a natural extension of $b_{n,k}$ and introduce its variant $\omega_{n,m,k}$ based on Theorem 2.1.
Let $n\geq m\geq k\geq 0$ and $(n,m,m)\vdash n+2m$. We place a vertical wall (bold red edge) between each pair of adjacent cells at the top row of the Young diagram of $(n,m,m)$. By removing any mkm-k cells at the top, we obtain a family of deformed Young diagrams with walls. One such diagram is illustrated in Fig. 5 for $n=9$, $m=7$, and $k=4$. Similarly, a deformed Young tableau with walls of this type is a filling of the cells by labels $1,2,\dots,n+m+k$, such that each label appears exactly once and the labels are increasing along each row and column, but two labels separated by a wall need not be increasing. Let $b_{n,m,k}$ denote the total number of deformed Young tableaux with walls on this family of deformed Young diagrams with walls.
Figure 5: A case in a family of deformed Young diagrams with walls for $n=9$, $m=7$, and $k=4$.
When $k=0$, the hook-length formula [23, Corollary 7.21.6] for counting standard Young tableau gives
$b_{n,m,0}=\frac{(n+m)!(n-m+1)}{m!(n+1)!}.$
Let
$C(x)=\sum_{n\geq 0}\mathrm{Cat}(n)x^{n}=\sum_{n\geq 0}\frac{1}{n+1}\binom{2n}{n}x^{n}=\frac{1-\sqrt{1-4x}}{2x}$
be the generating function of Catalan numbers. Then we obtain
$\sum_{0\leq m\leq n}b_{n,m,0}x^{n}y^{m}=\frac{1-yC(xy)}{1-x-y}.$
Proposition 2.4.
For $0\leq k\leq n$, we have $b_{n,n,k}=b_{n,k}$.
Proof.
A simple bijection readily yields $b_{n,n,k}=b_{n,k}$. For example, a Young tableau with walls in the count of $b_{n,n,k}$ can be transformed into one in the count of $b_{n,k}$ by first flipping it vertically, then horizontally, and finally replacing each entry $i$ with 2n+k+1i2n+k+1-i. ∎
Proposition 2.5.
The sequence $b_{n,m,k}$ is uniquely determined by the recurrence relation
$b_{n,m,k}=(m-k+1)b_{n,m,k-1}+b_{n,m-1,k}+b_{n-1,m,k}$ (10)
for $n\geq 2$ and $0\leq k\leq m\leq n$ with the initial conditions
$b_{0,0,0}=b_{1,0,0}=b_{1,1,0}=b_{1,1,1}=1\quad\text{and}\quad b_{m-1,m,k}=b_{n,k-1,k}=0.$
Proof.
The first part of the initial condition clearly holds. The equalities $b_{m-1,m,k}=b_{n,k-1,k}=0$ are related with the recurrence
$b_{n,m,k}=(m-k+1)\cdot b_{n,m,k-1}+b_{n,m-1,k}+b_{n-1,m,k}\quad\text{ for }\quad n\geq 1,\quad 0\leq k\leq m\leq n,$
which is obtained by considering where the largest element $v=n+m+k$ is located. This is classified by the following three cases:
Case 1: $v$ is in the bottom row, which forces $n\geq m+1$. Removing $v$ gives a Young tableau with walls counted by $b_{n-1,m,k}$, but note that $b_{m-1,m,k}$ has to be interpreted as 0;
Case 2: $v$ is in the middle row, which forces $m\geq k+1$. Removing $v$ gives a Young tableau with walls counted by $b_{n,m-1,k}$, but note that $b_{n,k-1,k}$ has to be interpreted as 0;
Case 3: $v$ is in the top row. Removing $v$ gives a Young tableau with walls counted by $b_{n,m,k-1}$, while inserting $v$ back have m(k1)m-(k-1) choices.
This completes the proof. ∎
Inspired by Theorem 2.1 and Proposition 2.5, we introduce the following variant of $b_{n,m,k}$.
Definition 2.6 (A variant of $b_{n,m,k}$).
We define the sequence $\omega_{n,m,k}$ by the recurrence
$\omega_{n,m,k}=\omega_{n-1,m+1,k}-(m-k+2)\omega_{n-1,m+1,k-1}-\omega_{n-2,m+1,k}$ (11)
for $n+m\geq 2$ and $0\leq k-1\leq m$, with the initial conditions $\omega_{1,0,0}=1$, $\omega_{-1,m,k}=0$ and
$\omega_{0,m,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot\frac{2^{m-k}}{(m-k+1)!}\cdot(2m+k+i-1)!!,$ (12)
where $\gamma_{j}$ satisfy recurrence relation (8) and initial condition $\gamma_{0}=1$.
The rest of this section is devoted to proving the following relationship, which implies Theorem 1.1 in view of Theorem 2.1.
Theorem 2.7.
We have $\omega_{n,m,k}=b_{n+m,m,k}$.
2.3 Two key lemmas
The most technical part of the proof of Theorem 2.7 lies in showing that $\omega_{n,k-1,k}=0$. For this purpose, we establish two auxiliary lemmas, the first of which is an explicit recursion for $\omega_{n,k-1,k}$.
Lemma 2.8.
For any $1\leq s\leq n$, the sequence $\omega_{n,k-1,k}$ satisfies the following recursion:
$\omega_{n,k-1,k}=\sum_{p=1}^{\left\lfloor\frac{s+1}{2}\right\rfloor}\sum_{q=1}^{s+2-2p}\alpha_{s}(p,q)\cdot\omega_{n-s-1,k+s-p,k+1-q}+\sum_{p=1}^{\left\lceil\frac{s+1}{2}\right\rceil}\sum_{q=1}^{s+3-2p}-\alpha_{s+1}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q},$
where
$\alpha_{s}(p,q)=\frac{(-1)^{q-p+1}\cdot(s-1+q-p)!}{(s-q-2p+2)!\cdot(q-1)!\cdot 2^{q-1}\cdot(p-1)!}.$ (13)
Proof.
The proof is by induction on $s$. If $s=1$, then we have
$\omega_{n,k-1,k}$ $=\alpha_{1}(1,1)\cdot\omega_{n-2,k,k}-\alpha_{2}(1,1)\omega_{n-1,k,k}-\alpha_{2}(1,2)\omega_{n-1,k,k-1}$
$=-\omega_{n-2,k,k}+\omega_{n-1,k,k}-\omega_{n-1,k,k-1}.$
This is consistent with the recurrence for $\omega_{n,k-1,k}$ in (11), hence the lemma holds for $s=1$. Assuming the formula to hold for s1s-1, we will prove it for $s$.
By the recurrence for $\omega_{n,m,k}$ in (11), we now have
$\omega_{n,k-1,k}$ $=\sum_{p=1}^{\left\lfloor\frac{s}{2}\right\rfloor}\sum_{q=1}^{s+1-2p}\alpha_{s-1}(p,q)\cdot\omega_{n-s,k+s-1-p,k+1-q}+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\omega_{n-s+1,k+s-1-p,k+1-q}$
$=\sum_{p=1}^{\left\lfloor\frac{s}{2}\right\rfloor}\sum_{q=1}^{s+1-2p}\alpha_{s-1}(p,q)\cdot\omega_{n-s,k+s-1-p,k+1-q}+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}$
$+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\big(-(s+q-p)\omega_{n-s,k+s-p,k-q}-\omega_{n-s-1,k+s-p,k+1-q}\big)$
$=\sum_{p=1}^{\left\lfloor\frac{s+1}{2}\right\rfloor}\sum_{q=1}^{s+2-2p}\alpha_{s}(p,q)\cdot\omega_{n-s-1,k+s-p,k+1-q}+\sum_{p=1}^{\left\lfloor\frac{s}{2}\right\rfloor}\sum_{q=1}^{s+1-2p}\alpha_{s-1}(p,q)\cdot\omega_{n-s,k+s-1-p,k+1-q}$
$+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}\alpha_{s}(p,q)\cdot(s+q-p)\omega_{n-s,k+s-p,k-q},$
where we use the fact $\lceil\frac{s}{2}\rceil=\lfloor\frac{s+1}{2}\rfloor$ in the last equality. It suffices to prove the following identity:
$\sum_{p=2}^{\left\lfloor\frac{s}{2}\right\rfloor+1}\sum_{q=1}^{s+3-2p}\alpha_{s-1}(p-1,q)\cdot\omega_{n-s,k+s-p,k+1-q}+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}$
$\quad+\sum_{p=1}^{\left\lceil\frac{s}{2}\right\rceil}\sum_{q=2}^{s+3-2p}\alpha_{s}(p,q-1)\cdot(s+q-p-1)\omega_{n-s,k+s-p,k+1-q}$ (14)
$=$ $\sum_{p=1}^{\left\lceil\frac{s+1}{2}\right\rceil}\sum_{q=1}^{s+3-2p}-\alpha_{s+1}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}.$
We begin by verifying the case $p=1$, that is, Eq. (14) reduces to
$\sum_{q=1}^{s}-\alpha_{s}(1,q)\omega_{n-s,k+s-1,k+1-q}+\sum_{q=2}^{s+1}\alpha_{s}(1,q-1)(s+q-2)\omega_{n-s,k+s-1,k+1-q}$
$=$ $\sum_{q=1}^{s+1}-\alpha_{s+1}(1,q)\omega_{n-s,k+s-1,k+1-q}.$
This is divided into the following three subcases:
  1. 1.
    When $q=1$, the equality $\alpha_{s}(1,1)=-\alpha_{s+1}(1,1)$ follows from (13).
  2. 2.
    When $q=s+1$, the identity $\alpha_{s}(1,s)\cdot(2s-1)=-\alpha_{s+1}(1,s+1)$ holds by (13).
  3. 3.
    When $2\leq q\leq s$, the equality $\alpha_{s}(1,q)+\alpha_{s}(1,q-1)\cdot(s+q-2)=-\alpha_{s+1}(1,q)$ still follows from (13).
We now turn to the case $p>1$. This requires us to distinguish the parity of $s$. We begin by assuming $s$ is even. When $p$ attains its maximum value of $\frac{s+2}{2}$, $q$ must be $1$, and we obtain $\alpha_{s-1}(\frac{s}{2},1)=-\alpha_{s+1}(\frac{s+2}{2},1)$ by (13). When $2\leq p\leq\frac{s}{2}$, we need only prove
$\sum_{q=1}^{s+3-2p}\alpha_{s-1}(p-1,q)\cdot\omega_{n-s,k+s-p,k+1-q}+\sum_{q=1}^{s+2-2p}-\alpha_{s}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}$
$\quad+\sum_{q=2}^{s+3-2p}\alpha_{s}(p,q-1)\cdot(s+q-p-1)\omega_{n-s,k+s-p,k+1-q}$ (15)
$=$ $\sum_{q=1}^{s+3-2p}-\alpha_{s+1}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}.$
This is also divided into the following three subcases:
  1. 1.
    When $q=1$, the equality $\alpha_{s-1}(p-1,1)-\alpha_{s}(p,1)=-\alpha_{s+1}(p,1)$ follows from (13).
  2. 2.
    When q=s+32pq=s+3-2p, the identity α_s-1(p-1,s+3-2p) +α_s(p,s+2-2p)⋅(2s-3p+2)=-α_s+1(p,s+3-2p) holds by (13).
  3. 3.
    When $2\leq q\leq s+2-2p$, the equality α_s-1(p-1,q)-α_s(p,q)+ α_s(p,q-1)⋅(s+q-p-1)=-α_s+1(p,q) still follows from (13).
Next, we consider the case of odd $s$. When $p$ attains its maximum value of $\frac{s+1}{2}$, we need only prove
$\sum_{q=1}^{2}\alpha_{s-1}(p-1,q)\cdot\omega_{n-s,k+s-p,k+1-q}+\sum_{q=1}^{1}-\alpha_{s}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}$
$\quad+\sum_{q=2}^{2}\alpha_{s}(p,q-1)\cdot(s+q-p-1)\omega_{n-s,k+s-p,k+1-q}$
$=$ $\sum_{q=1}^{2}-\alpha_{s+1}(p,q)\cdot\omega_{n-s,k+s-p,k+1-q}.$
At this time, when $q=1$, the identity $\alpha_{s-1}(\frac{s+1}{2}-1,1)-\alpha_{s}(\frac{s+1}{2},1)=-\alpha_{s+1}(\frac{s+1}{2},1)$ follows from (13); when $q=2$, the identity $\alpha_{s-1}(\frac{s+1}{2}-1,2)+\alpha_{s}(\frac{s+1}{2},1)\cdot\frac{s+1}{2}=-\alpha_{s+1}(\frac{s+1}{2},2)$. Similarly, when $2\leq p\leq\frac{s-1}{2}$, we need prove (15). The proof follows the same argument as for even $s$ and is omitted.
We have now proven that (14) holds, and that the formula of $\omega_{n,k-1,k}$ also holds for $s$. This completes the proof by induction. ∎
Next, we need the following two constant term facts to prove our second lemma.
Let $\mathrm{CT}_{x}F(x)$ denote the constant term of the formal Laurent series expansion of $F(x)$.
Fact 1.
Let $i$ be a nonnegative integer. If $k$ is a nonnegative integer, possibly less than $i$, or if $k$ is a half integer, then
$\mathop{\mathrm{CT}}_{x}\frac{(1+x)^{k}}{x^{i}}=\binom{k}{i}$.
Fact 2.
If $F(x)$ only contains nonnegative power of $x$, then
$\mathop{\mathrm{CT}}_{x}F(x)\frac{1}{1-u/x}=\mathop{\mathrm{CT}}_{x}F(x)\sum_{n\geq 0}\left(\frac{u}{x}\right)^{n}=F(u),$
where $u$ is independent of $x$.
This fact follows easily by linearity of the constant term operator. For rigorous definitions, see [24].
Lemma 2.9.
For any $n,k,i\in\mathbb{N}$ with $i\leq k$, we have
$\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}\frac{(-1)^{q-p}\cdot 2^{n-p}}{(n+3-2p-q)!\cdot(p-1)!}\cdot\binom{k-i}{q-1}\cdot\frac{(2n+4k-2p-2q+1-i)!!}{(4k-3-i)!!}$
$=$ $2^{n-1}\binom{n+3k-2}{n}$.
Proof.
We denote the left-hand side of the equation by $\mathrm{LHS}$. We obtain
$\mathrm{LHS}$ $=\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}(-1)^{p-q}2^{n-p}\binom{n+2-p-q}{p-1}\binom{k-i}{q-1}\cdot$
$\quad\quad\quad\cdot\frac{(4k-i-1)(4k-i+1)\cdots(4k-i+1+2n-2p-2q)}{(n+2-p-q)!}$
$=\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}(-1)^{p-q}2^{2n-2p-q+2}\binom{n+2-p-q}{p-1}\binom{k-i}{q-1}\cdot$
$\quad\quad\quad\cdot\binom{2k+n-p-q+\frac{1-i}{2}}{n+2-p-q}$.
By Fact 1, our problem thus reduces to computing the constant term in the following equation:
$\mathrm{LHS}$ $=\mathop{\mathrm{CT}}_{x,y,z}\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}(-1)^{p-q}2^{2n-2p-q+2}\frac{(1+x)^{k-i}}{x^{q-1}}\cdot\frac{(1+y)^{n+2-p-q}}{y^{p-1}}\cdot$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\cdot\frac{(1+z)^{2k+n-p-q+\frac{1-i}{2}}}{z^{n+2-p-q}}.$
At this time, extending the upper bounds of the summations is equivalent to adding some zeros by Fact 1. Consequently, we obtain
$\mathrm{LHS}$ $=\mathop{\mathrm{CT}}_{x,y,z}\sum_{p=1}^{\infty}\sum_{q=1}^{\infty}(-1)^{p-q}2^{2n-2p-q+2}\frac{(1+x)^{k-i}(1+y)^{n+2-p-q}(1+z)^{2k+n-p-q+\frac{1-i}{2}}}{x^{q-1}y^{p-1}z^{n+2-p-q}}.$
Applying the geometric series formula yields
$\mathrm{LHS}$ $=\mathop{\mathrm{CT}}_{x,y,z}\sum_{p=1}^{\infty}\frac{(-1)^{p+1}2^{2n-2p+1}(1+y)^{n+1-p}}{(1+x)^{i-k}y^{p-1}z^{n+1-p}(1+z)^{p-2k-n+\frac{i+1}{2}}}\cdot\frac{1}{1+\frac{z}{2x(1+y)(1+z)}}$
$=\mathop{\mathrm{CT}}_{x,y,z}\frac{2^{2n-1}(1+y)^{n}}{(1+x)^{i-k}z^{n}(1+z)^{-2k-n+\frac{i+3}{2}}}\cdot\frac{1}{1+\frac{z}{4y(1+z)(1+y)}}\cdot\frac{1}{1+\frac{z}{2x(1+y)(1+z)}}.$
We first eliminate $x$ using Fact 2, where $u=\frac{-z}{2(1+y)(1+z)}$. This gives
$\mathrm{LHS}$ $=\mathop{\mathrm{CT}}_{y,z}\frac{2^{2n-1}(1+y)^{n}}{z^{n}(1+z)^{-2k-n+\frac{i+3}{2}}\left(1+\frac{z}{4y(1+z)(1+y)}\right)\left(1-\frac{z}{2(1+y)(1+z)}\right)^{i-k}}.$
Next, we eliminate $y$. Only the denominator factor $1+\frac{z}{4y(1+y)(1+z)}$ may results in negative powers of $y$. By quadratic formula, we can write
$\frac{1}{1+\frac{z}{4y(1+y)(1+z)}}=\frac{y(1+y)}{(y-Y_{1})(y-Y_{2})}=\frac{(1+y)}{(1-Y_{1}/y)(y-Y_{2})},$
where
$Y_{1}=\frac{1-\sqrt{1+z}}{2\sqrt{1+z}}=O(z)\quad\text{and}\quad Y_{2}=\frac{-1-\sqrt{1+z}}{2\sqrt{1+z}}=-1+O(z).$
Then apply Fact 2 with $u=Y_{1}$ gives
$\mathrm{LHS}=\mathop{\mathrm{CT}}_{z}2^{n-2}z^{-n}(1+z)^{n+2k-\frac{i+3}{2}}\frac{(1+\sqrt{1+z})^{n+1}}{(\sqrt{1+z})^{n}}\cdot\left(\frac{1+z+\sqrt{1+z}}{(1+z)(1+\sqrt{1+z})}\right)^{k-i}.$
Note that some of the above computations were performed via Maple [17]. According to
$\frac{1+z+\sqrt{1+z}}{(1+z)(1+\sqrt{1+z})}=\frac{\sqrt{1+z}}{1+z}=\frac{1}{\sqrt{1+z}},$
we obtain
$\mathrm{LHS}$ $=\mathop{\mathrm{CT}}_{z}2^{n-2}z^{-n}(1+z)^{n+2k-\frac{i+3}{2}}\frac{(1+\sqrt{1+z})^{n+1}}{(\sqrt{1+z})^{n}}\cdot\left(\frac{1}{\sqrt{1+z}}\right)^{k-i}$
$=2^{n-2}\cdot\mathop{\mathrm{CT}}_{z}\frac{(1+z)^{\frac{3k+n-3}{2}}(1+\sqrt{1+z})^{n+1}}{z^{n}}.$
It is noteworthy that the above expression is independent of $i$. We assume the reader is familiar with the knowledge of residues. Informally, the residue can be viewed as the coefficient of the $z^{-1}$ term in a Laurent series expansion. Therefore, we have
$\mathrm{LHS}$ $=2^{n-2}\cdot\mathop{\mathrm{Res}}_{z=0}\frac{(1+z)^{\frac{3k+n-3}{2}}(1+\sqrt{1+z})^{n+1}}{z^{n+1}}$
$=2^{n-2}\cdot\frac{1}{2\pi i}\oint_{|z|=\varepsilon}\frac{(1+z)^{\frac{3k+n-3}{2}}(1+\sqrt{1+z})^{n+1}}{z^{n+1}}dz,$
where the integration path is taken near $z=0$, with $\varepsilon$ sufficiently small. Let $z=4t+4t^{2}$. Then we have $dz=4(1+2t)dt$ and
$\mathrm{LHS}$ $=2^{n-2}\cdot\frac{1}{2\pi i}\oint_{|t|=\varepsilon}\frac{(1+2t)^{3k+n-3}(2(1+t)^{n+1})}{(4t(1+t))^{n+1}}\cdot 4(1+2t)dt$
$=2^{n-2}\cdot\frac{1}{2\pi i}\oint_{|t|=\varepsilon}\frac{(1+2t)^{3k+n-2}}{2^{n-1}t^{n+1}}dt$
$=2^{-1}\cdot\mathop{\mathrm{Res}}_{t=0}\frac{(1+2t)^{3k+n-2}}{t^{n+1}}=2^{-1}\cdot\mathop{\mathrm{CT}}_{t}\frac{(1+2t)^{3k+n-2}}{t^{n}}$
$=2^{n-1}\binom{3k+n-2}{n}$.
This completes the proof. ∎
2.4 Finishing the proof of Theorem 1.1
In this subsection, we prove Theorem 2.7, which completes the proof of Theorem 1.1.
Lemma 2.10.
For any $n,k\in\mathbb{N}$, we have $\omega_{n,k-1,k}=0$.
Proof.
Taking $s=n$ in Lemma 2.8 gives
$\omega_{n,k-1,k}=\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{n+3-2p}-\alpha_{n+1}(p,q)\cdot\omega_{0,k+n-p,k+1-q}.$ (16)
By Eqs. (13) and (12), we have
$\omega_{n,k-1,k}$
$=$ $\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{n+3-2p}\frac{(-1)^{q-p}\cdot 2^{n-p}}{(n+3-2p-q)!\cdot(q-1)!\cdot(p-1)!}\sum_{i=0}^{k+1-q}\frac{\gamma_{k+1-q-i}}{i!}\cdot(2n+3k-2p-q+i)!!$
$=$ $\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{n+3-2p}\frac{(-1)^{q-p}\cdot 2^{n-p}}{(n+3-2p-q)!\cdot(q-1)!\cdot(p-1)!}\sum_{i=0}^{k+1-q}\frac{\gamma_{i}}{(k+1-q-i)!}\cdot(2n+4k-2p-2q+1-i)!!,$
where $\gamma_{j}$ satisfy recurrence relation (8) and initial condition $\gamma_{0}=1$. By interchanging the order of the summation above, we obtain
$\omega_{n,k-1,k}$ $=\sum_{i=0}^{k}\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}\frac{(-1)^{q-p}\cdot 2^{n-p}}{(n+3-2p-q)!\cdot(q-1)!\cdot(p-1)!}\cdot$
$\quad\quad\quad\quad\quad\quad\cdot\frac{\gamma_{i}}{(k+1-q-i)!}\cdot(2n+4k-2p-2q+1-i)!!$
$=\sum_{i=0}^{k}\sum_{p=1}^{\left\lceil\frac{n+1}{2}\right\rceil}\sum_{q=1}^{\min\{n+3-2p,k+1-i\}}\frac{(-1)^{q-p}\cdot 2^{n-p}}{(n+3-2p-q)!\cdot(q-1)!\cdot(p-1)!}\cdot$
$\quad\cdot\frac{(k-i)!}{(k+1-q-i)!}\cdot\frac{(2n+4k-2p-2q+1-i)!!}{(4k-3-i)!!}\cdot\frac{\gamma_{i}}{(k-i)!}\cdot(4k-3-i)!!,$
By Lemma 2.9, the above expression reduces to
$\omega_{n,k-1,k}$ $=\sum_{i=0}^{k}2^{n-1}\binom{n+3k-2}{n}\cdot\frac{\gamma_{i}}{(k-i)!}\cdot(4k-3-i)!!,$
$=2^{n-1}\binom{n+3k-2}{n}\sum_{i=0}^{k}\frac{\gamma_{i}}{(k-i)!}\cdot(4k-3-i)!!$
$=0,$
where the last equality follows from (8). This completes the proof of the lemma. ∎
We are now ready for the proof of Theorem 2.7.
Proof of Theorem 2.7.
For convenience, we set $\tilde{\omega}_{n,m,k}=b_{n+m,m,k}$. By the recurrence (10) of $b_{n,m,k}$, we have
$b_{n+m,m,k}=(m-k+1)b_{n+m,m,k-1}+b_{n+m,m-1,k}+b_{n+m-1,m,k}.$
Thus, we have the following recurrence relation for
$\tilde{\omega}_{n,m,k}=(m-k+1)\tilde{\omega}_{n,m,k-1}+\tilde{\omega}_{n+1,m-1,k}+\tilde{\omega}_{n-1,m,k}.$ (17)
The initial conditions for $b_{n,m,k}$ in Proposition 2.5 gives
$\tilde{\omega}_{0,0,0}=\tilde{\omega}_{1,0,0}=\tilde{\omega}_{0,1,0}=\tilde{\omega}_{0,1,1}=1,\,\,\tilde{\omega}_{-1,m,k}=b_{m-1,m,k}=0$ (18)
and
$\tilde{\omega}_{n,k-1,k}=b_{n+k-1,k-1,k}=0.$ (19)
Note that recurrence (17) for $\tilde{\omega}_{n,m,k}$ together with the initial conditions in (18) and (19) determines $\tilde{\omega}_{n,m,k}$ uniquely. To finish the proof, it remains to show that $\omega_{n,m,k}$ satisfies the same recurrence relation and initial conditions as that of $\tilde{\omega}_{n,m,k}$.
Setting $n\leftarrow n+1$ in (11) yields
$\omega_{n+1,m,k}=\omega_{n,m+1,k}-(m-k+2)\omega_{n,m+1,k-1}-\omega_{n-1,m+1,k},$
which can be recast as
$\omega_{n,m+1,k}=(m-k+2)\omega_{n,m+1,k-1}+\omega_{n+1,m,k}+\omega_{n-1,m+1,k}.$
Subsequently, setting $m\leftarrow m-1$ leads to
$\omega_{n,m,k}=(m-k+1)\omega_{n,m,k-1}+\omega_{n+1,m-1,k}+\omega_{n-1,m,k}.$
This shows that $\omega_{n,m,k}$ satisfies the same recurrence relation as that of $\tilde{\omega}_{n,m,k}$ in (17). Lemma 2.10 ensures that $\omega_{n,k-1,k}=0=\tilde{\omega}_{n,k-1,k}$. Finally, verifying the other initial conditions is straightforward:
$\omega_{0,0,0}=\omega_{1,0,0}=\omega_{0,1,0}=\omega_{0,1,1}=1,\,\,\omega_{-1,m,k}=0.$
This completes the proof of the theorem. ∎
Combining Proposition 2.4 with Theorem 2.7 yields $b_{n,k}=b_{n,n,k}=\omega_{0,n,k}$, which results in the following expression for $b_{n,k}$ in view of (12).
Theorem 2.11.
Let $0\leq k\leq n$. We have
$b_{n,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot\frac{2^{n-k}}{(n-k+1)!}\cdot(2n+k+i-1)!!,$
where $\gamma_{j}$ satisfy recurrence relation (8) and initial condition $\gamma_{0}=1$.
Proof of Theorem 1.1.
The proof now follows from Theorem 2.11 and Theorem 2.1. ∎
Corollary 2.12.
The $b_{n,k}$ satisfies the following recurrence relation:
$b_{n,k}=\frac{n-k+2}{2}\cdot b_{n,k-1}+\frac{2(2n+k-1)}{n-k+1}\cdot b_{n-1,k}.$
Proof.
This follows from Theorem 1.1 and the recurrence (2). ∎
2.5 Enumeration of tree-child networks
We now present several corollaries concerning the enumeration of tree-child networks. Pons and Batle [18, Proposition 8-12] derived the following results, Corollaries 2.13, 2.14, 2.15, 2.16, and 2.17 under the assumption that Conjecture 1.2 holds. We can now remove this assumption.
Corollary 2.13.
The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
$(n-k)|\mathcal{TC}_{n,k}|=(n+1-k)(n-k)|\mathcal{TC}_{n,k-1}|+n(2n+k-3)|\mathcal{TC}_{n-1,k}|,$
where initial values $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
Proof.
By combining Equations (6) and (2), we obtain this result. ∎
Corollary 2.14.
The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
$(n-k)!\cdot|\mathcal{TC}_{n,k}|=\sum_{i=0}^{k}n(2n+i-3)(n-1-i)!\cdot|\mathcal{TC}_{n-1,i}|,$
where initial values $|\mathcal{TC}_{1,0}|=1$ and $|\mathcal{TC}_{i,-1}|=|\mathcal{TC}_{i,i}|=0$ for any $i$.
Proof.
By recurrence (2), we obtain
$a_{n,k+1}=\sum_{i=k+1}^{n}\prod_{j=i+1}^{n}(2j+k)a_{i,k}.$
The result follows from the relation above combined with Eq (6). ∎
Corollary 2.15.
The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
$|\mathcal{TC}_{k+m+1,k}|=\sum_{\ell=0}^{m}(\ell+2)\left[\prod_{i=\ell+1}^{m}\left(1+\frac{k}{i+1}\right)(2i+3k-1)\right]|\mathcal{TC}_{k+\ell+1,k-1}|$
for $k\geq 1$.
Proof.
By Eq. (9), we obtain
$a_{n+1,k}=\sum_{i=0}^{k}(2n+i+1)a_{n,i}.$
This result follows immediately from the above expression and Eq (6). ∎
Corollary 2.16.
Let $\delta_{j}=j!\cdot\gamma_{j}$ for $j\geq 0$. The cardinalities $|\mathcal{TC}_{n,k}|$ satisfy
$|\mathcal{TC}_{n,k}|=\binom{n}{k}\sum_{i=0}^{k}\binom{k}{i}(2n+2k-i-3)!!\cdot\delta_{i},$
where the coefficients $\delta_{i}$ are determined by the recursive equation
$\delta_{i}=-\sum_{j=1}^{i}\binom{i}{j}\frac{(2i+j-3)!!}{(3i-3)!!}\delta_{i-j},\quad\text{and}\quad\delta_{0}=1.$
Proof.
This is a consequence of Theorem 2.1 together with Eq (6). ∎
Pons and Batle [18, Proposition 7] determined the asymptotic behavior of $a_{n,k}$ as $n\rightarrow\infty$ with fixed $k$, and then, assuming Conjecture 1.2, derived the following result.
Corollary 2.17.
For $k$ fixed and $n\rightarrow\infty$, the cardinalities $|\mathcal{TC}_{n,k}|$ grow as
$|\mathcal{TC}_{n,k}|$ $=\binom{n}{k}\frac{\sqrt{2}}{e^{n}}(2n)^{n+k-1}\Bigg(1-\sqrt{\frac{\pi}{2}}k(2n)^{-\frac{1}{2}}+\frac{14k^{2}-26k+11}{12}(2n)^{-1}$
$-\sqrt{\frac{\pi}{2}}k\frac{31k^{2}-93k+70}{48}(2n)^{-\frac{3}{2}}+\frac{2900k^{4}-14376k^{3}+25264k^{2}-19332k+5565}{6048}(2n)^{-2}$
$\quad+O(n^{-\frac{5}{2}})\Bigg).$
3 Generating functions for $b_{n,k}$ and $b_{n,m,k}$
Introduce the generating functions for $b_{n,k}$ and $b_{n,m,k}$ as
$D_{k}(t)=\sum_{n\geq 0}b_{n,k}t^{n}\quad\text{and}\quad B_{k}(x,t)=\sum_{k\leq m\leq n}b_{n,m,k}x^{n-m}t^{m}.$
In this section, we obtain an explicit formula for $D_{k}(t)$ from Theorem 2.11 and derive a system of functional equations for $B_{k}(x,t)$ using the kernel method [3].
3.1 An explicit formula for $D_{k}(t)$
Theorem 3.1.
The generating function for $b_{n,k}$ is given by
$D_{k}(t)$ $=\sum_{j=0}^{\frac{k-1}{2}}\frac{\gamma_{k-2j-1}}{(2j+1)!\cdot 2^{j+\frac{3k+1}{2}}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot\binom{2j+3k-1}{j+\frac{3k-1}{2}}\cdot\frac{t^{k-1}}{(1-4t)^{j+\frac{3k}{2}}}$
$\quad\quad+\sum_{j=0}^{\frac{k-1}{2}}\frac{\gamma_{k-2j}\cdot 2^{j+\frac{3k-5}{2}}}{(2j)!}\cdot\left(j+\frac{3k-3}{2}\right)!\cdot\frac{t^{k-1}}{(1-4t)^{\frac{3k-1}{2}+j}}$
when $k$ is odd, and
$D_{k}(t)$ $=\sum_{j=0}^{\frac{k}{2}}\frac{\gamma_{k-2j}}{(2j)!\cdot 2^{j+\frac{3k}{2}}}\cdot\left(j+\frac{3k-2}{2}\right)!\cdot\binom{2j+3k-2}{j+\frac{3k-2}{2}}\cdot\frac{t^{k-1}}{(1-4t)^{j+\frac{3k-1}{2}}}$
$\quad\quad+\sum_{j=0}^{\frac{k}{2}-1}\frac{\gamma_{k-2j-1}\cdot 2^{j+\frac{3k-4}{2}}}{(2j+1)!}\cdot\left(j+\frac{3k-2}{2}\right)!\cdot\frac{t^{k-1}}{(1-4t)^{\frac{3k}{2}+j}}$
when $k$ is even.
Proof.
By Theorem 2.11, we have
$b_{n,k}=\sum_{i=0}^{k}\frac{\gamma_{k-i}}{i!}\cdot\frac{2^{n-k}}{(n-k+1)!}\cdot(2n+k+i-1)!!,$ (20)
where $\gamma_{j}$ satisfy the recursion (8) and the initial condition $\gamma_{0}=1$. We need to derive the formula for $D_{k}(t)$ according to the parity of $k$.
We begin with the case for $k$ being odd. It is convenience to rewrite (20) as
$b_{n,k}=\sum_{j=0}^{\left\lfloor\frac{k}{2}\right\rfloor}\frac{\gamma_{k-2j}}{(2j)!}\cdot\frac{(2n+k+2j-1)!!}{(n-k+1)!}2^{n-k}+\sum_{j=0}^{\left\lceil\frac{k}{2}\right\rceil-1}\frac{\gamma_{k-2j-1}}{(2j+1)!}\cdot\frac{(2n+k+2j)!!}{(n-k+1)!}2^{n-k}.$ (21)
When $k$ is odd and $0\leq j\leq\frac{k-1}{2}$, we first consider the following generating function:
$T_{1}:=\sum_{n\geq 0}\frac{(2n+k+2j)!!}{(n-k+1)!}\cdot 2^{n-k}t^{n}=\sum_{n\geq 0}\frac{2^{-j-\frac{3k+1}{2}}\cdot(2n+k+2j+1)!}{(n+j+\frac{k+1}{2})!\cdot(n-k+1)!}t^{n}.$
Setting $m=n+j+\frac{k+1}{2}$ gives
$T_{1}$ $=\sum_{m\geq j+\frac{k+1}{2}}\frac{2^{-j-\frac{3k+1}{2}}\cdot(2m)!}{m!\cdot(m-(j+\frac{3k-1}{2}))!}t^{m-j-\frac{k+1}{2}}$
$=2^{-j-\frac{3k+1}{2}}\cdot t^{-j-\frac{k+1}{2}}\sum_{m\geq j+\frac{k+1}{2}}\binom{2m}{m}\binom{m}{j+\frac{3k-1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{m}.$
By introducing a variable $z$ and extracting the coefficient of $z^{j+\frac{3k-1}{2}}$, we have
$T_{1}$ $=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}\sum_{m\geq j+\frac{k+1}{2}}\binom{2m}{m}[z^{j+\frac{3k-1}{2}}](1+z)^{m}t^{m}$
$=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}[z^{j+\frac{3k-1}{2}}]\sum_{m\geq 0}\binom{2m}{m}((1+z)t)^{m}.$
Since $\sum_{s\geq 0}\binom{2s}{s}t^{s}=\frac{1}{\sqrt{1-4t}}$, we obtain
$T_{1}$ $=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}[z^{j+\frac{3k-1}{2}}]\frac{1}{\sqrt{1-4(1+z)t}}$
$=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}[z^{j+\frac{3k-1}{2}}]\frac{1}{\sqrt{1-4t}}\cdot\frac{1}{\sqrt{1-\frac{4tz}{1-4t}}}.$
According to $(1-t)^{-\alpha}=1+\sum_{s\geq 1}\frac{\alpha(\alpha+1)\cdots(\alpha+s-1)}{s!}t^{s}$, we have
$T_{1}$ $=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}\frac{1}{\sqrt{1-4t}}[z^{j+\frac{3k-1}{2}}]\cdot\sum_{s\geq 0}\binom{2s}{s}\left(\frac{tz}{1-4t}\right)^{s}$
$=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot t^{-j-\frac{k+1}{2}}\frac{1}{\sqrt{1-4t}}\binom{2j+3k-1}{j+\frac{3k-1}{2}}\frac{t^{j+\frac{3k-1}{2}}}{(1-4t)^{j+\frac{3k-1}{2}}},$
which can be simplified to
$T_{1}=2^{-j-\frac{3k+1}{2}}\cdot\left(j+\frac{3k-1}{2}\right)!\cdot\binom{2j+3k-1}{j+\frac{3k-1}{2}}\frac{t^{k-1}}{(1-4t)^{j+\frac{3k}{2}}}.$ (22)
Secondly, we consider the following generating function:
$T_{2}:=\sum_{n\geq 0}\frac{(2n+k+2j-1)!!}{(n-k+1)!}\cdot 2^{n-k}t^{n}=\sum_{n\geq 0}\frac{2^{n+j+\frac{k-1}{2}}\cdot(n+j+\frac{k-1}{2})!\cdot 2^{n-k}}{(n-k+1)!}t^{n}.$
Setting m=nk+1m=n-k+1 and $k=2p+1$ (since $k$ is odd) yields
$T_{2}$ $=\sum_{m\geq 0}\frac{2^{m+j+3p}\cdot(m+j+3p)!\cdot 2^{m-1}}{m!}t^{m+2p}=2^{j+3p-1}t^{2p}\sum_{m\geq 0}\frac{(3p+m+j)!}{m!}(4t)^{m}$
$=2^{j+3p-1}t^{2p}\sum_{m\geq 0}\frac{(3p+j)!\cdot(3p+j+1)(3p+j+2)\cdots(3p+j+m)}{m!}(4t)^{m}$
$=2^{j+3p-1}(3p+j)!t^{2p}\cdot\frac{1}{(1-4t)^{3p+j+1}}.$
Therefore, we have
$T_{2}=2^{j+\frac{3k-5}{2}}\left(\frac{3k-3}{2}+j\right)!\frac{t^{k-1}}{(1-4t)^{\frac{3k-1}{2}+j}}.$ (23)
Combining (21), (22) and (23) establishes the desired formula for $D_{k}(t)$ for odd $k$.
Similarly, we can derive the generating function for $D_{k}(t)$ when $k$ is even. The details of the proof will be omitted due to the similarity. ∎
3.2 Functional equations for $B_{k}(x,t)$ via the kernel method
We aim to derive a system of functional equations for $B_{k}(x,t)$ using Proposition 2.5 and the kernel method. This is an independent approach for $D_{k}(t)$. Let $F_{0}(x,t)=1$ and
$F_{k}(x,t):=\sum_{k\leq m\leq n}(m-k+1)b_{n,m,k-1}x^{n-m}t^{m}$
for $k\geq 1$. It is clear that for $k\geq 1$,
$F_{k}(x,t)$ $=\sum_{0\leq k\leq m\leq n}m\cdot b_{n,m,k-1}x^{n-m}t^{m}+\sum_{0\leq k\leq m\leq n}(1-k)\cdot b_{n,m,k-1}x^{n-m}t^{m}$
$=t\cdot\frac{\partial B_{k-1}(x,t)}{\partial t}+(1-k)\cdot B_{k-1}(x,t).$
By the recurrence of $b_{n,m,k}$ in Proposition 2.5, we obtain
$B_{k}(x,t)$ $=F_{k}(x,t)+\sum_{0\leq k\leq m\leq n}b_{n,m-1,k}x^{n-m}t^{m}+\sum_{0\leq k\leq m\leq n}b_{n-1,m,k}x^{n-m}t^{m},$
$=F_{k}(x,t)+\frac{t}{x}\cdot B_{k}(x,t)-\frac{t}{x}\cdot D_{k}(t)+x\cdot B_{k}(x,t).$
Therefore, we have
$\left(1-x-\frac{t}{x}\right)B_{k}(x,t)=F_{k}(x,t)-\frac{t}{x}\cdot D_{k}(t).$ (24)
This is a special type of functional equation with two unknowns $B_{k}(x,t)$ and $D_{k}(t)$, where the factor $\left(1-x-\frac{t}{x}\right)$ on the left-hand side is called the kernel of (24). Solving the kernel $1-x-\frac{t}{x}=0$, i.e., $x^{2}-x+t=0$ for $x$ gives two solutions:
$X_{1}(t)=\frac{1+\sqrt{1-4t}}{2}=1+O(t)\quad\text{and}\quad X_{2}(t)=\frac{1-\sqrt{1-4t}}{2}=t\cdot C(t).$
According to the kernel method [3], we can substitute $x=X_{2}(t)$ to obtain
$D_{k}(t)=\left(\frac{x}{t}\cdot F_{k}(x,t)\right)\Big|_{x=X_{2}(t)}.$
The above functional equations are summarized in the following theorem.
Theorem 3.2.
For $k\geq 1$, we have the following system of functional equations
$\left\{\begin{aligned} &D_{k}(t)=\left(\frac{x}{t}\cdot F_{k}(x,t)\right)\Big|_{x=X_{2}(t)},\\ &F_{k}(x,t)=t\cdot\frac{\partial B_{k-1}(x,t)}{\partial t}+(1-k)\cdot B_{k-1}(x,t),\\ &B_{k}(x,t)=\frac{F_{k}(x,t)-\frac{t}{x}\cdot D_{k}(t)}{1-x-\frac{t}{x}},\end{aligned}\right$.
with the initial condition $F_{0}(x,t)=1$ and $D_{0}(t)=C(t)=\frac{1-\sqrt{1-4t}}{2t}$.
From a computational perspective, Theorem 3.2 is highly practical to compute expressions for $D_{k}(t)$ and $B_{k}(x,t)$. In the computational implementation, we perform the substitution t=u(1u)t=u(1-u), thus ensuring that we consistently work with rational functions and avoid fractional exponents. At this time, we have
$X_{2}(t)=u=\frac{1-\sqrt{1-4t}}{2},\quad B_{0}(x,t)=\frac{-1}{x-(1-u)}.$
and
$\frac{\partial B_{k-1}(x,t)}{\partial t}=\frac{1}{1-2u}\cdot\frac{\partial B_{k-1}(x,u(1-u))}{\partial u}.$
Using Theorem 3.2, we can easily compute expressions for $D_{k}(t)$ for $k\leq 140$ by Maple, which agree with the results from Theorem 3.1. This gives supports for the truth of Funchs and Yu’s enumerative conjecture (Theorem 1.1) in our earlier study.
4 Recurrence for $b_{n,k}$ via linear extensions of some posets
In our initial approach to Theorem 1.1, we began by computing the linear extensions of some posets associated with $b_{n,k}$. As a byproduct of this work, we provide formulas for the linear extensions of some posets. We ultimately obtain a monstrously large recurrence relation for $b_{n,k}$ as presented below.
Theorem 4.1.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$b_{n,k}=\binom{2n+k}{n}\cdot f_{n,k}-\sum_{j=1}^{n}\sum_{s=0}^{k}\sum_{m=0}^{s}\frac{(-1)^{2m}\cdot(j+k-s)\cdot(n-j-m)!\cdot(k+2j-m-1)!}{2^{k-s}\cdot(j-k+s)!\cdot(k-s)!\cdot j!\cdot(s-m)!\cdot(n-j-s)!}\cdot b_{n-j,m},$
where $f_{n,k}=\frac{(n-k+1)(n-k+2)\cdots(n+k)}{2^{k}\cdot k!}$.
The proof of Theorem 4.1 is given in appendix B; the outline of the proof steps are:
  1. 1.
    We present closed formulas for linear extensions of several poset families related to $b_{n,k}$; see Lemmas B.7, B.8, B.9.
  2. 2.
    We study a counting function $u_{n,k}$ associated with $b_{n,k}$, as defined in Section B.3. We derive two transformation formulas between $b_{n,k}$ and $u_{n,k}$; see Theorems B.10 and B.12. This is necessary to derive the recurrence relation for $b_{n,k}$.
  3. 3.
    By applying the inclusion-exclusion principle together with the previous lemmas, we obtain a recurrence relation for $b_{n,k}$; see Theorems B.14 and 4.1.
It is our expectation that the approach to counting standard Young tableaux with walls (or phylogenetic networks) via linear extensions and the principle of inclusion-exclusion will find applications in future research.
5 Concluding remarks, open problems
This paper investigates enumeration formulas for two specific types of Young tableaux with walls. The main achievement is the proof of Theorem 1.1, which leads to the verification of Pons and Batle’s enumerative conjecture on counting phylogenetic tree-child networks (see Conjecture 1.2). The enumeration of Young tableaux with walls is generally a challenging problem. This paper provides several methods and techniques to facilitate future related studies.
For future research, we highlight the following two open problems.
Problem 5.1.
Are there combinatorial proofs of Theorem 1.1, Conjecture 1.2 and Corollary 2.12?
Problem 5.2.
Can Theorem 4.1 be applied to prove Theorem 1.1 or Corollary 2.12?
Acknowledgments
We are grateful to Wenjie Fang for communicating Theorem 1.1 (Fuchs and Yu’s enumerative conjecture) to us during his visit to Shandong University in the summer of 2025, to Shu Xiao Li for insightful discussions on Proposition 2.5, and to Michael Fuchs and Guan-Ru Yu for kindly sharing the intriguing background information, as well as for their inspiring comments, regarding Theorem 1.1. This work was supported by the National Science Foundation of China (grants 12571355, 12322115 & 12271301) and the Fundamental Research Funds for the Central Universities.
Appendix A Appendix: Formulae for $a_{n,k}$ and $b_{n,k}$ for some fixed small $k$
We present formulas for $a(n,k)$ and $b_{n,k}$ here for observation of their structures.
We derived the following formulas for $a_{n,k}$ directly from recursion (9).
Proposition A.1.
Let $0\leq k\leq n$. We derive the formulas of $a_{n,k}$ for some small $k$:
$a_{n,0}=(2n-1)!!,$
$a_{n,1}=(2n+1)!!-(2n)!!,$
$a_{n,2}=\left(n+\frac{5}{3}\right)\cdot(2n+1)!!-(2n+2)!!,$
$a_{n,3}=\frac{n+3}{3}\cdot(2n+3)!!-\left(n+\frac{79}{48}\right)\cdot(2n+2)!!,$
$a_{n,4}=\left(\frac{n^{2}}{6}+\frac{7n}{6}+\frac{319}{189}\right)\cdot(2n+3)!!-\frac{(16n+31)(n+2)}{24}\cdot(2n+2)!!,$
$a_{n,5}=\left(\frac{63n^{2}+609n+1006}{1890}\right)(2n+5)\cdot(2n+3)!!-\left(\frac{1}{6}n^{2}+\frac{13}{16}n+\frac{9107}{9216}\right)\cdot(2n+4)!!.$
We derived the following formulas for $b_{n,k}$ directly from the recursion in Theorem 4.1.
Proposition A.2.
Let $0\leq k\leq n$. We derive the formulas of $b_{n,k}$ for some small $k$:
$b_{n,0}=\frac{1}{n+1}\binom{2n}{n}$,
$b_{n,1}=\frac{1}{2}\cdot\frac{(2n+1)!}{(n!)^{2}}-2^{2n-1},$
$b_{n,2}=\frac{n(n+\frac{5}{3})}{4}\cdot\frac{(2n+1)!}{(n!)^{2}}-n(n+1)\cdot 2^{2n-1},$
$b_{n,3}=\frac{(n+3)}{3\cdot 2^{4}}\cdot\frac{(2n+3)!}{(n-2)!(n+1)!}-(n-1)n(n+1)\left(n+\frac{79}{48}\right)\cdot 2^{2n-2},$
$b_{n,4}=\left(\frac{n^{2}}{6}+\frac{7n}{6}+\frac{319}{189}\right)\cdot\frac{(2n+3)!}{2^{5}\cdot(n-3)!(n+1)!}$
$\quad\quad\quad\quad\quad\quad-\frac{(16n+31)}{3}\cdot(n-2)(n-1)n(n+1)(n+2)\cdot 2^{2n-6}.$
$b_{n,5}=\left(\frac{1}{15}n^{3}+\frac{73}{90}n^{2}+\frac{5057}{1890}n+\frac{503}{189}\right)\cdot\frac{(2n+4)!}{2^{7}\cdot(n-4)!(n+2)!}$
$\quad\quad\quad-\left(\frac{1}{6}n^{2}+\frac{13}{16}n+\frac{9107}{9216}\right)\cdot(n-3)(n-2)(n-1)n(n+1)(n+2)\cdot 2^{2n-3}.$
Appendix B Appendix: Linear extensions of some posets
In this Appendix, we provide formulas for the linear extensions of some posets. Subsec. B.1 introduces the theory of $P$-partitions and linear extensions of posets, which is required later. Subsec. B.2 presents formulas for linear extensions of several poset families. This provides some necessary lemmas for deriving the recurrence formula for $b_{n,k}$. The relationship between $b_{n,k}$ and $u_{n,k}$ is characterized in Subsec. B.3. In Subsec. B.4, we establish a recurrence relation for $b_{n,k}$.
B.1 $P$-partitions and linear extensions
In this subsection, we introduce some concepts and useful lemmas regarding $P$-partitions and linear extensions.
Let $(P,\preceq)$ be a finite partially ordered set (or poset). We say that $t$ covers $s$ if $s\prec t$ and there is no $c\in P$ such that $s\prec c\prec t$. The Hasse diagram of a finite poset $P$ is a graph whose vertices are the elements of $P$, whose edges represent cover relations, and in which, if $s\prec t$, then $t$ is drawn above $s$ in the diagram.
A $P$-partition is a map $\sigma:P\rightarrow\mathbb{N}$ satisfying the following condition: If $s\prec t$ in $P$, then $\sigma(s)\geq\sigma(t)$; that is, $\sigma$ is order-reversing. In addition, if $\sum_{s\in P}\sigma(s)=n$, we say $\sigma$ is a $P$-partition of $n$ and write $|\sigma|=n$. It is often convenient to represent the relation $s\prec t$ by an arrow $s\to t$. In this notation, a $P$-partition corresponds to a vertex labeling $\sigma$ such that for every arrow $s\to t$, we have $\sigma(s)\geq\sigma(t)$.
The concept of $P$-partitions was first investigated by MacMahon [16, Subsections 439,441]. In 1970, Knuth [13] provided clarification on MacMahon’s work within the theory of $P$-partitions. Later, Knuth [13] offered a clarification of MacMahon’s results in his 1970 work on the theory of $P$-partitions. The first systematic treatment of this subject, however, was established by Stanley in his seminal papers [20, 21].
We denote the set of all $P$-partitions on the poset $P$ as $\pi(P)$. The fundamental generating function associated with $\pi(P)$ is defined as
$F_{P}(x_{1},x_{2},\ldots,x_{p})=\sum_{\sigma\in\pi(P)}x_{1}^{\sigma(1)}x_{2}^{\sigma(2)}\cdots x_{p}^{\sigma(p)}.$
By setting $x_{i}=q$ for all $i$, we obtain the enumerative generating function of $\pi(P)$:
$G_{P}(q)=F_{P}(q,q,\ldots,q)=\sum_{\sigma\in\pi(P)}q^{|\sigma|}.$
The set $p]:=\{1,2,\ldots,p\}$ with its usual order forms a $p$-element chain, that is, a poset in which all elements are totally ordered. This poset is denoted $\mathbf{p}$. Let #P=p\#P=p. An order-preserving bijection $\tau:P\rightarrow\mathbf{p}$ is called a linear extension of $P$. The number of linear extensions of $P$ is denoted by $e(P)$.
Now, let $(P,\preceq)$ be a finite poset on the set $[p]$. Any linear extension $\tau\colon P\to\mathbf{p}$ can be identified with the permutation $\tau^{-1}(1),\tau^{-1}(2),\dots,\tau^{-1}(p)$ of $[p]$. The collection of all such $e(P)$ permutations is denoted by $\mathcal{L}(P)$ and is called the Jordan–Hölder set of $P$. Counting linear extensions of a poset is a computationally difficult problem, and was shown by Brightwell and Winkler [5] to be #P-complete.
Let $\mu=\mu_{1}\mu_{2}\cdots\mu_{p}$ be a permutation of $[p]$. For $1\leq i\leq p-1$, if $\mu_{i}>\mu_{i+1}$, then $i$ is called a descent of $\mu$. The descent set of $\mu$ is defined as
$D_{\mu}=\{i\in[p-1]:\mu_{i}>\mu_{i+1}\}.$
Let $d(\mu)$ denote the number of descents of $\mu$, that is, $d(\mu)=\#D_{\mu}$.
Theorem B.1 ([22, Theorem 3.15.5]).
Let $(P,\preceq)$ be a finite poset on $p]:=\{1,2,\ldots,p\}$. Then
$F_{P}(x_{1},x_{2},\ldots,x_{p})=\sum_{\mu\in\mathcal{L}(P)}\frac{\prod_{j\in D_{\mu}}x_{\mu_{1}}x_{\mu_{2}}\cdots x_{\mu_{j}}}{\prod_{i=1}^{p}(1-x_{\mu_{1}}x_{\mu_{2}}\cdots x_{\mu_{i}})}.$
Define the major index $\mathrm{maj}(\mu)$ of $\mu$ by $\mathrm{maj}(\mu)=\sum_{j\in D_{\mu}}j$.
Theorem B.2 ([22, Theorem 3.15.7]).
Let $(P,\preceq)$ be a finite poset on $[p]$. Then
$G_{P}(q)=\frac{\sum_{\mu\in\mathcal{L}(P)}q^{\mathrm{maj}(\mu)}}{(1-q)(1-q^{2})\cdots(1-q^{p})}.$
The following result provides a formula for the number $e(P)$ of linear extensions of a poset.
Lemma B.3 ([22, Section 3]).
Let $(P,\preceq)$ be a finite poset on $[p]$. Then
$e(P)=p!\cdot\bigg(\Big((1-q)^{p}\cdot G_{P}(q)\Big)\Big|_{q=1}\bigg).$ (25)
The direct sum of $P_{1}$ and $P_{2}$ is the poset $P_{1}+P_{2}$ on the disjoint union $P_{1}\uplus P_{2}$ such that $s\preceq t$ in $P_{1}+P_{2}$ if (i) $s,t\in P_{1}$ and $s\preceq t$ in $P_{1}$, or (ii) $s,t\in P_{2}$ and $s\preceq t$ in $P_{2}$. The ordinal sum of posets $P_{1}$ and $P_{2}$, which is the poset $P_{1}\oplus P_{2}$ on the disjoint union $P_{1}\uplus P_{2}$ such that $s\preceq t$ in $P_{1}\oplus P_{2}$ if (i) $s,t\in P_{1}$ and $s\preceq t$ in $P_{1}$, or (ii) $s,t\in P_{2}$ and $s\preceq t$ in $P_{2}$, or (iii) $s\in P_{1}$ and $t\in P_{2}$.
Lemma B.4 ([22, Section 3.5]).
If $P=P_{1}+P_{2}+\cdots+P_{k}$ and $n_{i}=\#P_{i}$, then
$e(P)=\binom{n_{1}+n_{2}+\cdots+n_{k}}{n_{1},n_{2},\ldots,n_{k}}e(P_{1})e(P_{2})\cdots e(P_{k}).$
If $P=P_{1}\oplus P_{2}\oplus\cdots\oplus P_{k}$ and $n_{i}=\#P_{i}$, then
$e(P)=e(P_{1})e(P_{2})\cdots e(P_{k}).$
Let $I_{m}$ be the poset consisting of $m$ isolated vertices, i.e., an $m$-element antichain. Then we have $e(I_{m})=m!$. Let $L_{m}=I_{1}\oplus I_{1}\oplus\cdots\oplus I_{m}$ ($m$ times). Thus $L_{m}$ is an $m$-element chain. It follows that $e(L_{m})=1$.
Let $P$ be a poset with $p$ elements. In its Hasse diagram $\Gamma$, the relation $s\prec t$ is conventionally represented by an arrow $s\to t$. If there exists a directed path from vertex $v$ to vertex $u$ in $\Gamma$, we say that $v$ reaches $u$ and denote by $v\dashrightarrow u$. The number of vertices in $\Gamma$ that can reach vertex $u$ is denoted by $\omega(u)$.
When the Hasse diagram $\Gamma$ of the poset $P$ is a directed tree in which every vertex has out-degree either $1$ or 0, the enumerative generating function for $\pi(P)$ admits a simple form.
Lemma B.5 ([15]).
If the Hasse diagram $\Gamma$ of the poset $P$ is a directed tree in which every vertex has out-degree either $1$ or 0, then the generating function $G_{P}(q)$ is given by
$G_{P}(q)=\prod_{v\in\Gamma}\frac{1}{1-q^{\omega(v)}}.$
Furthermore, by Lemma B.3, we obtain
$e(P)=\frac{p!}{\prod_{v\in\Gamma}\omega(v)}.$
B.2 Linear extensions of a class of posets.
In this subsection, we present a closed formula for the linear extensions of a class of posets. This provides some necessary lemmas for deriving the recurrence formula for $b_{n,k}$.
Let $n$ be a positive integer. Given a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, define the poset $F_{(n;i_{1},i_{2},\ldots,i_{k})}$ as shown in Fig. 6, where the relation $s\prec t$ is conventionally represented by an arrow $s\to t$. We set $F_{(n;i_{1},i_{2},\ldots,i_{k})}=L_{n}$ ($n$-element chain) if $k=0$ by convention.
Figure 6: The Hasse diagram of the poset $F_{(n;i_{1},i_{2},\ldots,i_{k})}$.
We now define $f_{n,k}$ as the sum of the number of linear extensions of all posets $F_{(n;i_{1},i_{2},\ldots,i_{k})}$, i.e.,
$f_{n,k}:=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}e(F_{(n;i_{1},i_{2},\ldots,i_{k})}).$
Lemma B.6.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have $f_{n,0}=1$ and
$f_{n,k}=\sum_{i_{1}=1}^{n-k+1}\sum_{i_{2}=i_{1}+1}^{n-k+2}\cdots\sum_{i_{k}=i_{k-1}+1}^{n}i_{1}(i_{2}+1)(i_{3}+2)\cdots(i_{k}+k-1)\quad\text{for}\quad 1\leq k\leq n.$
Proof.
For a $n$-element chain, we know that $e(L_{n})=1$. It follows that $f_{n,0}=1$. Now we consider the cases $1\leq k\leq n$. Given a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, the Hasse diagram of the poset $F_{(n;i_{1},i_{2},\ldots,i_{k})}$ is a directed tree in which every vertex has out-degree either $1$ or 0. By definition of $\omega(v)$, we obtain
$\omega(v_{i_{1}})=i_{1}+1,\quad\omega(v_{i_{2}})=i_{2}+2,\quad\ldots,\quad\omega(v_{i_{k}})=i_{k}+k,$
and so
$\prod_{v\in F_{(n;i_{1},i_{2},\ldots,i_{k})}}\omega(v)=\frac{(n+k)!}{i_{1}(i_{2}+1)(i_{3}+2)\cdots(i_{k}+k-1)}.$
By Lemma B.5, we obtain
$e(F_{(n;i_{1},i_{2},\ldots,i_{k})})=\frac{(n+k)!}{(n+k)!}\cdot i_{1}(i_{2}+1)(i_{3}+2)\cdots(i_{k}+k-1).$
This completes the proof. ∎
Lemma B.7.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$f_{n,0}=1,\quad f_{n,k}=\frac{(n-k+1)(n-k+2)\cdots(n+k)}{2^{k}\cdot k!}\quad\text{for}\quad 1\leq k\leq n,$ (26)
and
$\sum_{n\geq 0}f_{n,k}x^{n}=\frac{(2k-1)!!\cdot x^{k}}{(1-x)^{2k+1}}.$ (27)
Proof.
We know that
$f_{n,k}=\sum_{i_{1}=1}^{n-k+1}\sum_{i_{2}=i_{1}+1}^{n-k+2}\cdots\sum_{i_{k}=i_{k-1}+1}^{n}i_{1}(i_{2}+1)(i_{3}+2)\cdots(i_{k}+k-1)\quad\text{for}\quad 1\leq k\leq n.$
The above equation can be written equivalently as
$f_{n,k}=\sum_{c_{1}\geq 1;\ c_{k}\leq n+k-1\atop c_{i}-c_{i-1}\geq 2\text{ for }2\leq i\leq k}c_{1}c_{2}\cdots c_{k}.$
We consider the following two cases according to $c_{k}$: 1): When $c_{k}=n+k-1$, we have $c_{k-1}\leq c_{k}-2=n+k-3$ and
$f_{n,k}$ $=(n+k-1)\sum_{c_{1}\geq 1;\ c_{k-1}\leq n+k-3\atop c_{i}-c_{i-1}\geq 2\text{ for }2\leq i\leq k-1}c_{1}c_{2}\cdots c_{k-1}$
$=(n+k-1)\cdot f_{n-1,k-1}.$
2): When $c_{k}\neq n+k-1$, we get $f_{n,k}=f_{n-1,k}$. Therefore, we obtain
$f_{n,k}=f_{n-1,k}+(n+k-1)\cdot f_{n-1,k-1}.$
Let
$g_{n,k}=\frac{(n-k+1)(n-k+2)\cdots(n+k)}{2^{k}\cdot k!}.$
One can verify that $g_{n,k}=g_{n-1,k}+(n+k-1)\cdot g_{n-1,k-1}$. It is clear that $f_{n,k}$ and $g_{n,k}$ satisfy the same initial conditions. It follows that
$f_{n,k}=\frac{(n-k+1)(n-k+2)\cdots(n+k)}{2^{k}\cdot k!}\quad\text{for}\quad 1\leq k\leq n.$
Now we have
$\sum_{n\geq 0}f_{n,k}x^{n}$ $=\sum_{n\geq k}\frac{(n-k+1)(n-k+2)\cdots(n+k)}{2^{k}\cdot k!}x^{n}$
$=\frac{(2k)!}{k!\cdot 2^{k}}\sum_{n\geq k}\frac{(n+k)(n+k-1)\cdots(n-k+1)}{(2k)!}x^{n}$
$=\frac{(2k)!}{k!\cdot 2^{k}}\sum_{n\geq k}\binom{n+k}{2k}x^{n}=\frac{(2k)!}{k!\cdot 2^{k}}\sum_{m\geq 0}\binom{2k+m}{m}x^{m+k}$
$=\frac{(2k-1)!!\cdot x^{k}}{(1-x)^{2k+1}}.$
This completes the proof. ∎
Note that the sequences $(f_{n,1})_{n\geq 1}$, $(f_{n,2})_{n\geq 2}$ and $(f_{n,3})_{n\geq 3}$ are the sequences [A240440], [A050534] and [A000217] in the OEIS [19], respectively.
We now consider the linear extensions of the direct sum of the poset $F_{(n;i_{1},i_{2},\ldots,i_{k})}$ and an $n$-element chain $L_{n}$.
Lemma B.8.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}e(F_{(n;i_{1},i_{2},\ldots,i_{k})}+L_{n})=\binom{2n+k}{n}\cdot f_{n,k}.$
Proof.
This result follows directly from Lemma B.4. ∎
Given a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, define the poset $\widetilde{F}_{(n;i_{1},i_{2},\ldots,i_{k})}$ as shown in Fig. 7, where the relation $s\prec t$ is conventionally represented by an arrow $s\to t$.
Figure 7: The Hasse diagram of the poset $\widetilde{F}_{(n;i_{1},i_{2},\ldots,i_{k})}$.
Similarly, we define $\widetilde{f}_{n,k}$ as the sum of the number of linear extensions of all posets $\widetilde{F}_{(n;i_{1},i_{2},\ldots,i_{k})}$, i.e.,
$\widetilde{f}_{n,k}=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}e(\widetilde{F}_{(n;i_{1},i_{2},\ldots,i_{k})}).$
Lemma B.9.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$\widetilde{f}_{n,k}=\binom{2n+k-1}{n}\cdot f_{n,k}.$
Proof.
It is clear that $\omega(v_{n})=2n+k$. When $k=0$, by Lemma B.5, we have
$\widetilde{f}_{n,0}=\frac{(2n)!}{n!(n-1)!(2n)}=\binom{2n-1}{n}$.
When $1\leq k\leq n$, by Lemma B.5, we obtain
$\widetilde{f}_{n,k}=\frac{(2n+k)!}{n!(n+k-1)!(2n+k)}\cdot\sum_{i_{1}=1}^{n-k+1}\sum_{i_{2}=i_{1}+1}^{n-k+2}\cdots\sum_{i_{k}=i_{k-1}+1}^{n}i_{1}(i_{2}+1)(i_{3}+2)\cdots(i_{k}+k-1).$
This completes the proof. ∎
B.3 Characterization of $b_{n,k}$
Let $n$ be a positive integer. Given a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, define the poset $D_{(n;i_{1},i_{2},\ldots,i_{k})}$ as shown in Fig. 8, where the relation $s\prec t$ is conventionally represented by an arrow $s\to t$.
Figure 8: The Hasse diagram of the poset $D_{(n;i_{1},i_{2},\ldots,i_{k})}$.
By the definition of $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ we see that this number is also equal to the number of labelings of the poset with integers $1,2,\ldots,2n+k$, such that every row and column is increasing along the direction of the arrow. Therefore, we obtain
$b_{n,k}=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}e(D_{(n;i_{1},i_{2},\ldots,i_{k})}).$
Similarly, define the poset $U_{(n;i_{1},i_{2},\ldots,i_{k})}$ as shown in Fig. 9, where the relation $s\prec t$ is conventionally represented by an arrow $s\to t$. In other words, $U_{(n;i_{1},i_{2},\ldots,i_{k})}$ is obtained by reversing $k$ downward arrows in $D_{(n;i_{1},i_{2},\ldots,i_{k})}$ to upward arrows.
Figure 9: The Hasse diagram of the poset $U_{(n;i_{1},i_{2},\ldots,i_{k})}$.
We now define $u_{n,k}$ as the sum of the number of linear extensions of all posets $U_{(n;i_{1},i_{2},\ldots,i_{k})}$, i.e.,
$u_{n,k}=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}e(U_{(n;i_{1},i_{2},\ldots,i_{k})}).$
We now derive a transformation formula between $b_{n,k}$ and $u_{n,k}$.
Theorem B.10.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$b_{n,k}=\sum_{i=0}^{k}(-1)^{i}\binom{2n+k}{k-i}\binom{n-i}{k-i}\cdot(k-i)!\cdot u_{n,i}.$
Proof.
This formula is derived by applying the inclusion–exclusion principle and Lemma B.4. We first illustrate the above formula with an example before proving the general case.
Let $n=4$ and $k=2$. Given $i_{1}=2,i_{2}=3$, we obtain the decomposition via the inclusion-exclusion principle as shown in Fig. 10.
Figure 10: An example for the decomposition via the inclusion-exclusion principle.
The equation in Fig. 10 means that
$e(D_{(4;2,3)})=e(P_{1})-(e(P_{2})+e(P_{3}))+e(P_{4}).$
Recall that $I_{m}$ denotes an $m$-element antichain. It is clear that
$e(P_{1})=e(I_{2}+U_{(4;0)})=\binom{10}{2}\cdot 2!\cdot\mathrm{Cat}(4)$
and
$e(D_{(4;2,3)})=e(I_{2}+U_{(4;0)})-(e(I_{1}+U_{(4;2)})+e(I_{1}+U_{(4;3)}))+e(U_{(4;2,3)}).$
The above discussion can be generalized to the general case. Given a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, by an application of the inclusion–exclusion principle, we obtain
$e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ $=\sum_{j=0}^{k}(-1)^{j}\sum_{(h_{1},\ldots,h_{j})\subseteq(i_{1},\ldots,i_{k})}e(I_{k-j}+U_{(n;h_{1},\ldots,h_{j})})$
$=\sum_{j=0}^{k}(-1)^{j}\sum_{(h_{1},\ldots,h_{j})\subseteq(i_{1},\ldots,i_{k})}\binom{2n+k}{k-j}e(I_{k-j})e(U_{(n;h_{1},\ldots,h_{j})})$
$=\sum_{j=0}^{k}(-1)^{j}\binom{2n+k}{k-j}(k-j)!\sum_{(h_{1},\ldots,h_{j})\subseteq(i_{1},\ldots,i_{k})}e(U_{(n;h_{1},\ldots,h_{j})}).$
The summation of both sides over all integer sequences $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$ yields
$b_{n,k}$ $=\sum_{j=0}^{k}(-1)^{j}\binom{2n+k}{k-j}(k-j)!\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}\sum_{(h_{1},\ldots,h_{j})\subseteq(i_{1},\ldots,i_{k})}e(U_{(n;h_{1},\ldots,h_{j})})$
$=\sum_{j=0}^{k}(-1)^{j}\binom{2n+k}{k-j}(k-j)!\sum_{1\leq h_{1}<\cdots<h_{j}\leq n}\binom{n-j}{k-j}e(U_{(n;h_{1},\ldots,h_{j})})$
$=\sum_{j=0}^{k}(-1)^{j}\binom{2n+k}{k-j}(k-j)!\binom{n-j}{k-j}u_{n,j}.$
This completes the proof. ∎
By Theorem B.10, we get the following result.
Corollary B.11.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$(-1)^{k}u_{n,k}=b_{n,k}-\sum_{i=0}^{k-1}(-1)^{i}\binom{2n+k}{k-i}\binom{n-i}{k-i}\cdot(k-i)!\cdot u_{n,i}.$
A proof similar to that of Theorem B.10 gives the following result.
Theorem B.12.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$u_{n,k}=\sum_{i=0}^{k}(-1)^{i}\binom{2n+k}{k-i}\binom{n-i}{k-i}\cdot(k-i)!\cdot b_{n,i}.$
B.4 Recurrence for $b_{n,k}$
Let $n$ be a positive integer. Given a positive integer sequence
$1\leq i_{1}<\cdots<i_{k-s}\leq j<i_{k-s+1}<\cdots<i_{k}\leq n,$
define the poset $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$ as shown in Fig. 11, where the relation $s\prec t$ is conventionally represented by an arrow $s\to t$. In Fig. 11, among the downward arrows from the upper layer, there are $s$ to the right of vertex $v_{j}$ and ksk-s to the left.
From an alternative perspective, the Hasse diagram of the poset $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$ is formed by connecting those of posets $\widetilde{F}_{(j;i_{1},i_{2},\ldots,i_{k-s})}$ and $D_{(n-j;i_{k-s+1}-j,i_{k-s+2}-j,\ldots,i_{k}-j)}$ with a directed edge $v_{j}\to v_{j+1}$. For convenience, we denote $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$ by
$\widetilde{F}_{(j;i_{1},i_{2},\ldots,i_{k-s})}\rightarrow D_{(n-j;i_{k-s+1}-j,i_{k-s+2}-j,\ldots,i_{k}-j)}$
Figure 11: The Hasse diagram of the poset $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$.
We define $r_{n,k}$ as the sum of the number of linear extensions of all posets $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$, i.e.,
$r_{n,k}=\sum_{j=1}^{n}\sum_{s=0}^{k}\sum_{1\leq i_{1}<\cdots<i_{k-s}\leq j\atop j<i_{k-s+1}<\cdots<i_{k}\leq n}e(R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}).$
Lemma B.13.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$r_{n,k}=\sum_{j=1}^{n}\sum_{s=0}^{k}\binom{2j+k-s-1}{j}\cdot f_{j,k-s}\cdot\left(\sum_{i=0}^{s}(-1)^{i}\binom{2n+k}{s-i}\binom{n-j-i}{s-i}(s-i)!\cdot u_{n-j,i}\right).$
Proof.
We know that the Hasse diagram of the poset $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$ is formed by connecting those of posets $\widetilde{F}_{(j;i_{1},i_{2},\ldots,i_{k-s})}$ and $D_{(n-j;i_{k-s+1}-j,i_{k-s+2}-j,\ldots,i_{k}-j)}$ with a directed edge $v_{j}\to v_{j+1}$. By Theorem B.10, the Hasse diagram of the poset $D_{(n-j;i_{k-s+1}-j,i_{k-s+2}-j,\ldots,i_{k}-j)}$ can be decomposed using the inclusion-exclusion principle. Similarly, for the Hasse diagram of poset $R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}$, we can retain the $\widetilde{F}_{(j;i_{1},i_{2},\ldots,i_{k-s})}$ portion and apply the same decomposition to its latter part. For each decomposed Hasse diagram, its linear extensions can be computed by employing Lemma B.4.
For fixed $s$ and $j$, Lemma B.9 and Theorem B.10 yield
$\sum_{1\leq i_{1}<\cdots<i_{k-s}\leq j<i_{k-s+1}<\cdots<i_{k}\leq n}e(R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})})$
$=$ $\left(\binom{2j+k-s-1}{j}\cdot f_{j,k-s}\right)\cdot\left(\sum_{i=0}^{s}(-1)^{i}\binom{2n+k}{s-i}\binom{n-j-i}{s-i}(s-i)!\cdot u_{n-j,i}\right).$
This completes the proof. ∎
Theorem B.14.
Let $n$ be a positive integer and $0\leq k\leq n$. Then we have
$b_{n,k}$ $=\binom{2n+k}{n}\cdot f_{n,k}$
$\quad-\sum_{j=1}^{n}\sum_{s=0}^{k}\binom{2j+k-s-1}{j}\cdot f_{j,k-s}\cdot\left(\sum_{i=0}^{s}(-1)^{i}\binom{2n+k}{s-i}\binom{n-j-i}{s-i}(s-i)!\cdot u_{n-j,i}\right).$
Proof.
By the definition of $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ we see that the number $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ is equal to the number of labelings of the poset with entries the distinct integers $1,2,\ldots,2n+k$, such that every row and column is increasing along the direction of the arrow.
For a positive integer sequence $1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n$, we observe the Hasse diagram of the poset $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$: two rightward lines are connected by $n$ downward arrows. Thus, the number $e(D_{(n;i_{1},i_{2},\ldots,i_{k})})$ equals the number $e(F_{(n;i_{1},i_{2},\ldots,i_{k})}+L_{n})$ minus the number of linear extensions of all posets containing an upward arrow. These posets with up arrows can be partitioned into $n$ classes according to the position of the rightmost up arrow, where each class is illustrated in Fig. 11 for a fixed $j$. Therefore, we obtain
$b_{n,k}=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}e(F_{(n;i_{1},i_{2},\ldots,i_{k})}+L_{n})-\sum_{j=1}^{n}\sum_{s=0}^{k}\sum_{1\leq i_{1}<\cdots<i_{k-s}\leq j\atop j<i_{k-s+1}<\cdots<i_{k}\leq n}e(R_{(n;i_{1},\ldots,i_{k-s},j,i_{k-s+1},\ldots,i_{k})}).$
The theorem follows from Lemmas B.8 and B.13. ∎
Now, we can show Theorem 4.1 by using Lemma B.7, Theorem B.12 and Theorem B.14.
Proof of Theorem 4.1.
According to Lemma B.7, Theorem B.12, and Theorem B.14, it follows that
$b_{n,k}$ $=\binom{2n+k}{n}\cdot f_{n,k}-\sum_{j=1}^{n}\sum_{s=0}^{k}\binom{2j+k-s-1}{j}\cdot f_{j,k-s}\cdot$
$\quad\quad\quad\quad\cdot\Bigg(\sum_{i=0}^{s}(-1)^{i}\binom{2n+k}{s-i}\binom{n-j-i}{s-i}(s-i)!\cdot$
$\quad\quad\quad\quad\quad\quad\quad\quad\cdot\sum_{m=0}^{i}(-1)^{m}\binom{2(n-j)+i}{i-m}\binom{n-j-m}{i-m}(i-m)!\cdot b_{n-j,m}\Bigg),$
where $f_{n,k}$ has closed formula given in (26). Therefore, we obtain
$b_{n,k}$ $=\binom{2n+k}{n}\cdot f_{n,k}-\sum_{j=1}^{n}\sum_{s=0}^{k}\sum_{i=0}^{s}\sum_{m=0}^{i}\binom{2j+k-s-1}{j}\cdot f_{j,k-s}(-1)^{i+m}\cdot$
$\qquad\cdot\Bigg(\binom{2n+k}{s-i}\binom{n-j-i}{s-i}(s-i)!\cdot\binom{2(n-j)+i}{i-m}\binom{n-j-m}{i-m}(i-m)!\cdot b_{n-j,m}\Bigg)$
$=\binom{2n+k}{n}\cdot f_{n,k}-\sum_{j=1}^{n}\sum_{s=0}^{k}\sum_{m=0}^{s}\frac{(-1)^{2m}\cdot(j+k-s)\cdot(n-j-m)!\cdot(k+2j-m-1)!}{2^{k-s}\cdot(j-k+s)!\cdot(k-s)!\cdot j!\cdot(s-m)!\cdot(n-j-s)!}\cdot b_{n-j,m},$
where the last equality is obtained from Maple. ∎